if p arithmetic means are inserted between 5 and 41 so that the ratio A3/Ap-1= 2/5 . Find p

Grade:11

1 Answers

Himanshu Rathour
5 years ago
If p terms are there between 5 and 41 the common difference d=41-5 =36/p+1
                         p+1
A3=5+3(36/p+1)
Ap-1=5+(p-1)(36/p+1)
    A3         =     5p+5+108            =  2
   Ap-1            5p+5+36+36p           5
p=9